# A Joint Representation Using Continuous and Discrete Features for Cardiovascular Diseases Risk Prediction on Chest CT Scans

### v1

@@ Introduction
Cardiovascular diseases (CVD) continue to pose a significant public health concern, accounting for 45% of deaths in Europe [wilkins2017european] and 31% in the United States [mozaffarian2016heart]. Despite a notable 7% reduction in CVD mortality rates since 2005, largely attributed to advances in clinical screening and prevention strategies, heart disease remains one of the leading causes of death in the United States [d20191997]. The ongoing challenge is to identify asymptomatic patients who would benefit most from preventive measures, such as initiating statin therapy [ridker2008rosuvastatin] and others, in a convenient, accessible, accurate, and non-invasive manner. This is a crucial aspect of effective preventive cardiology patient management.
@@
Several guidelines endorse different CVD risk prediction models. The 2010 guidelines [greenland20102010] from the American College of Cardiology / American Heart Association (ACC / AHA) suggest the use of the Framingham Risk Score (FRS) [wilson1998prediction, eichler2007prediction, d2008general], a well-established and thoroughly validated model that considers traditional risk factors such as cholesterol levels and blood pressure, etc. In contrast, the 2016 European guidelines favor the Systematic Coronary Risk Evaluation (SCORE) model [piepoli2016guidelines]. The 2016 Chinese guidelines introduce the China-PAR model, specifically designed to estimate the 10-year risk for atherosclerotic cardiovascular disease [yang2016predicting, liu2018predicting]. However, these recommended risk prediction models rely on a limited set of predictors, which may result in suboptimal risk management performances on various patient populations.
@@
Recent studies have begun to utilize CT scans to extract quantitative imaging biomarkers, demonstrating good potential for surpassing the predictive performance of multivariate FRS [pickhardt2020automated]. With the developments of deep learning in medical imaging, previous work explored integrating CT scans with deep learning algorithms to automatically obtain CT quantitative biomarkers related to CVD risk [xu2023ai, eng2021automated, zeleznik2021deep]. The methods specialized in multivariate analysis are designed for CVD risk prediction and can contain considerable interpretability. However, they may often overlook the interactions among different factors, and due to restrictive modeling assumptions and a limited number of predictors, these methods tend to attain only moderate predictive performance [siontis2012comparisons]. On the other hand, some studies [van2019direct, chao2021deep] have directly applied deep learning to predict CVD risk from chest CT scans. While these methods can greatly improve the predictive capacity and accuracy, they provide little interpretable insight into the model’s decision-making process.
@@
In this study, we propose a simple yet effective information fusion method, called DeepCVD, that extracts continuous deep features and calculates discrete quantitative biomarkers from CT scans. These features are subsequently integrated using a joint representation module that fosters interaction, ultimately enhancing the precision of CVD risk prediction and enabling the analysis of each feature’s contribution to the model’s decision-making process. In the feature extraction phase, discrete features are generated from image segmentation models as CT quantitative biomarkers, crafted based on current clinical knowledge to improve the model’s predictive accuracy and applicability. Meanwhile, continuous deep image features derived from a trained deep CVD classifier encompass rich higher-order semantic contextual information pertinent to CVD, albeit challenging to quantify. We propose a joint representation module to fully utilize the complementary information of these two channels of features for superior CVD risk prediction. Specifically, we employ an instance-wise feature gating mechanism (IFGM) to align the densities and dimensions of the input continuous and discrete features. We apply a Gated Residual Network (GRN) [lim2021temporal], a flexible and non-linear operation, to obtain more robust feature embeddings. Inspired by the multi-head attention mechanism [vaswani2017attention], we have designed a soft instance-wise feature interaction mechanism (SIFIM), facilitating thorough interactions among features across diverse subspaces and capturing their intricate interrelations for joint representation learning. Importantly, this strategy preserves the autonomy of each input feature, enabling physicians to understand their respective contributions to the model’s decisions easily. Furthermore, our feature interaction technique allows the model to adaptively output the contribution of each feature for various CVD, aligning closely with practical clinical settings. Our approach has achieved competitive performance using both public datasets and a private external validation patient cohort.
@@ Results
The LDCT-NLST cohort is a significant dataset intended to evaluate the effectiveness of low-dose chest CT scans in lung cancer screening. In this trial, it also collects information related to the CVD of subjects, making the dataset applicable for research in CVD risk prediction [chao2021deep, van2019direct]. Each subject undergoes one to three CT examinations, each producing multiple CT scans employing different CT reconstruction kernels. It has 33,413 CT volumes from 10,395 subjects, and each subject is labeled as either CVD-Positive or CVD-Negative. A subject is considered CVD positive if any cardiovascular abnormalities are reported in their CT screening examinations, or if the subject dies of CVD. CVD-negative subjects have no history of CVD during the clinical trial, no cardiovascular abnormalities reported in any CT scans, and do not die from circulatory system diseases.
@@
We developed the continuous and discrete features joint representation CVD risk prediction system (DeepCVD) in the LDCT-NLST cohort (Fig. 1a). We adopt the three random subsets generated by the prior study [chao2021deep] entirely, with the training set accounting for 70% (7,268 subjects), the validation set 10% (1,042 subjects), and the independent test set 20% (2,085 subjects). We found that the LDCT-NLST cohort contained some data that did not meet the requirements of our model, such as incomplete field of view (FOV), missing slices, and severe cardiac artifacts. Two radiologists cleansed the dataset and removed the problematic subjects, resulting in 10,325 subjects. The training set comprises 7,227 subjects, the validation set comprises 1,033 subjects, and the test set includes 2,065 subjects. Our goal was to achieve accurate CVD risk prediction using chest CT scans while providing physicians with reliable bases for model decisions.
@@
The proposed DeepCVD was used to identify subjects with CVDs in the LDCT-NLST testing cohort, which consisted of 2,065 patients (583 patients CVD-Positive and 1,482 patients CVD-Negative). The quantitative performance was summarized in Table 1, and the receiver operating characteristic curves (ROCs) of multiple methods were shown in Fig. 1c. DeepCVD achieved an area under the curve (AUC) of 0.875 (95% Confidence Interval (CI), 0.857-0.891), a sensitivity of 0.616 (95% CI, 0.576-0.655), and a specificity of 0.918 (95% CI, 0.903-0.931). From Table 1, it was evident that models using our designed quantitative biomarkers or directly using prior knowledge, such as pericoronary calcification and epicardial fat, whether Xgboost [chen2016Xgboost], Tri2D-Net [chao2021deep], or our proposed DeepCVD, demonstrated significantly better specificity compared to models like ResNet34 [he2016deep] that did not use prior knowledge. However, the complex encoding of prior knowledge can impair the ability to suppress false positives (non-encoded Xgboost specificity is 0.971, while encoded Tri2D-Net and DeepCVD had specificities of 0.952 and 0.918, respectively). By combining quantitative biomarkers and continuous deep features, DeepCVD increased the sensitivity by 27.0% and the F1 score by 11.9% compared to Tri2D-Net, at the expense of slightly reduced specificity.
To further present the quality of the embeddings of different methods in the LDCT-NLST testing cohort, t-SNE [van2008visualizing] was used to visualize the embeddings of different methods. In Fig. 2, the color of the nodes corresponded to CVD-Positive and CVD-Negative, verifying the discriminative power of the method. As indicated in Fig. 2, DeepCVD and Tri2D-Net can achieve more compact and separated clusters compared with other methods that do not incorporate prior insights.
@@
To assess the generalizability of our proposed DeepCVD, we directly applied the model trained on the LDCT-NLST dataset to the NERC-MBD standard-dose chest CT dataset. DeepCVD achieved an AUC of 0.843 (95% CI, 0.837-0.849), accuracy of 0.819 (95% CI, 0.813-0.824), sensitivity of 0.756 (95% CI, 0.746-0.765), and specificity of 0.870 (95% CI, 0.863-0.876). These results showed significant performance improvement (p ) compared to the previous state-of-the-art approach Tri2D-Net [chao2021deep] (accuracy increased by 9.8%, sensitivity by 28.4%, F1 score by 17.1%, and AUC by 5.9%, while the specificity remained almost the same). In comparison, other methods had inferior performance that was statistically significant (all p ). Detailed evaluation performance was summarized in Table 2.
@@ Discussion
This study thoroughly validates the effectiveness of DeepCVD. It successfully combines the advantages of discrete and continuous features, improves CVD risk prediction performance, and demonstrates strong generalizability. Multisource feature fusion techniques are also widely used in the medical domain, as feature concatenation is a commonly adopted and effective method for joint feature representation [kiela2014learning]. In the first stage, we concatenate continuous and discrete features in this manner and then feed them into the second stage features joint representation module. We observe an increase in AUC from 0.844 to 0.862 compared to directly using a CVD risk classifier with ResNet34. However, this mode of feature interaction does not fully leverage the respective advantages of quantitative discrete biomarkers and deep continuous features, with specificity increasing from 0.896 to 0.956, while sensitivity decreased from 0.543 to 0.440 (as indicated in Table 1 and Fig. 6c). Furthermore, although our model is trained on LDCT-NLST, it demonstrates remarkable generalization to standard-dose chest CT scans as shown in Table 2. Moreover, our joint representation module outperformed models like Xgboost, even when only discrete features were used, demonstrating the benefits of effective feature interaction for enhancing model performance (shown in Fig. 6d).
@@
We have also discovered that incorporating prior knowledge, either through direct predictions based on CT quantitative biomarkers using Xgboost or by integrating them into deep models such as Tri2D-Net, effectively improves the model’s specificity. Tri2D-Net incorporated pericardial fat and calcification as strong constraints in the prediction model. In the real world, not all CVDs are strongly associated with just two biomarkers, which reduces the model’s sensitivity and its ability to generalize to real-world scenarios. As shown in Fig. 2, although Tri2D-Net also demonstrates high-quality embeddings across the LDCT-NLST test set, its performance significantly drops on the external test set (see Table 2). This is mainly because the external test set includes many subjects with cerebral infarction (refer to Extended Data Fig. 3). Consequently, using CAM [selvaraju2017grad] for model interpretability also becomes less meaningful. In contrast, our model aligns and flexibly processes discrete and continuous features through the instance-wise feature-gated mechanism to generate more robust embeddings. Then, the soft instance-wise feature interaction mechanism achieves thorough interaction and fusion of features while maintaining relative independence. This ensures that our model optimally balances specificity and sensitivity (as seen in Table 1 and Table 2), and automatically learns the relationships between different CVDs and various biomarkers and deep features (as seen in Fig. 3). This informs physicians about the role of each feature in the decision-making process and opens up possibilities to discover new clinical biomarkers in the CVD domain.
@@
A recent study utilized a mature and fully automated abdominal CT-based algorithm with predefined metrics to quantify aortic calcification, muscle density, the ratio of visceral to subcutaneous fat, liver fat, and bone mineral density for assessing CVD risk [pickhardt2020automated]. The research demonstrated that the multivariate combination of CT biomarkers could effectively enhance CVD risk prediction performance over traditional risk factors. For example, the combination of four CT-based quantitative biomarkers—aortic calcification, muscle density, the ratio of visceral to subcutaneous fat, and liver fat—resulted in a 2-year AUC of 0.817 (95% CI 0.768-0.866). Although this study used a different retrospective cohort and abdominal non-contrast CT, the results of the two studies both demonstrated the potential CVD risk prediction value of harnessing the rich biometric tissue data embedded within all body CT scans that typically go unused in routine practice. However, this approach requires predefined CT biomarkers, and the combination of these biomarkers to enhance CVD risk prediction performance needs to be validated through repeated experiments. Our approach goes a step further by using deep features to represent those CT biomarkers that cannot or have not yet been predefined, thereby enhancing CVD risk prediction performance. Furthermore, through a unique design, our model can output the contribution of each feature in the decision-making process. This not only provides more insights for doctors but also suppresses those CT biomarkers that do not aid in CVD risk prediction.
@@
In clinical practice, the fusion of imaging information with clinical information can result in increased accuracy, mode informative clinical decision-making, and improved patient outcomes [huang2020fusion]. Our DeepCVD can seamlessly integrate into clinical information. In future work, we aim to expand the discrete features from CT quantitative biomarkers to multimodal biomarkers, incorporating laboratory indicators closely linked to CVDs such as blood pressure, cholesterol, and smoking history into the model to enhance predictive performance. Additionally, we plan to establish connections between biomarkers and specific CVDs through contribution scores, providing actionable guidance for physicians in subsequent diagnoses and achieving greater clinical value.
@@ Methods
This study aims to accurately predict CVD risk using chest CT scans and provide physicians with reliable information for model decisions. Our method is based on two findings and assumptions. First, while well-known quantitative biomarkers from CT scans, such as the coronary artery calcium score, are considered risk factors for CVD [pickhardt2020automated, iacobellis2022epicardial], we believe that many undefined or difficult-to-quantify biomarkers related to CVD have not been fully exploited within chest CT scans. Fully utilizing these biomarkers can effectively improve the performance of CVD risk prediction. Second, the relevance and contributions of biomarkers should vary across different CVDs, requiring these features to interact sufficiently to address the diversity of CVDs while also maintaining relative independence so that the model can adaptively output the contribution of each biomarker in decision-making.

### v2

@@ Introduction
Cardiovascular diseases (CVD) continue to pose a significant public health concern, accounting for 45% of deaths in Europe [1] and 31% in the United States [2]. Despite a notable 7% reduction in CVD mortality rates since 2005, largely attributed to advances in clinical screening and prevention strategies, heart disease remains one of the leading causes of death in the United States [3]. The ongoing challenge is to identify asymptomatic patients who would benefit most from preventive measures, such as initiating statin therapy [4] and others, in a convenient, accessible, accurate, and non-invasive manner. This is a crucial aspect of effective preventive cardiology patient management.
@@
Several guidelines endorse different CVD risk prediction models. The 2010 guidelines [5] from the American College of Cardiology / American Heart Association (ACC / AHA) suggest the use of the Framingham Risk Score (FRS) [6, 7, 8], a well-established and thoroughly validated model that considers traditional risk factors such as cholesterol levels and blood pressure, etc. In contrast, the 2016 European guidelines favor the Systematic Coronary Risk Evaluation (SCORE) model [9]. The 2016 Chinese guidelines introduce the China-PAR model, specifically designed to estimate the 10-year risk for atherosclerotic cardiovascular disease [10, 11]. However, these recommended risk prediction models rely on a limited set of predictors, which may result in suboptimal risk management performances on various patient populations.
@@
Recent studies have begun to utilize CT scans to extract quantitative imaging biomarkers, demonstrating good potential for surpassing the predictive performance of multivariate FRS [12]. With the developments of deep learning in medical imaging, previous work explored integrating CT scans with deep learning algorithms to automatically obtain CT quantitative biomarkers related to CVD risk [13, 14, 15]. The methods specialized in multivariate analysis are designed for CVD risk prediction and can contain considerable interpretability. However, they may often overlook the interactions among different factors, and due to restrictive modeling assumptions and a limited number of predictors, these methods tend to attain only moderate predictive performance [16]. On the other hand, some studies [17, 18] have directly applied deep learning to predict CVD risk from chest CT scans. While these methods can greatly improve the predictive capacity and accuracy, they provide little interpretable insight into the model’s decision-making process.
@@
In this study, we propose a simple yet effective information fusion method, called DeepCVD, that extracts continuous deep features and calculates discrete quantitative biomarkers from CT scans. These features are subsequently integrated using a joint representation module that fosters interaction, ultimately enhancing the precision of CVD risk prediction and enabling the analysis of each feature’s contribution to the model’s decision-making process. In the feature extraction phase, discrete features are generated from image segmentation models as CT quantitative biomarkers, crafted based on current clinical knowledge to improve the model’s predictive accuracy and applicability. Meanwhile, continuous deep image features derived from a trained deep CVD classifier encompass rich higher-order semantic contextual information pertinent to CVD, albeit challenging to quantify. We propose a joint representation module to fully utilize the complementary information of these two channels of features for superior CVD risk prediction. Specifically, we employ an instance-wise feature gating mechanism (IFGM) to align the densities and dimensions of the input continuous and discrete features. We apply a Gated Residual Network (GRN) [19], a flexible and non-linear operation, to obtain more robust feature embeddings. Inspired by the multi-head attention mechanism [20], we have designed a soft instance-wise feature interaction mechanism (SIFIM), facilitating thorough interactions among features across diverse subspaces and capturing their intricate interrelations for joint representation learning. Importantly, this strategy preserves the autonomy of each input feature, enabling physicians to understand their respective contributions to the model’s decisions easily. Furthermore, our feature interaction technique allows the model to adaptively output the contribution of each feature for various CVD, aligning closely with practical clinical settings. Our approach has achieved competitive performance using both public datasets and a private external validation patient cohort.
@@ Results
The LDCT-NLST cohort is a significant dataset intended to evaluate the effectiveness of low-dose chest CT scans in lung cancer screening. In this trial, it also collects information related to the CVD of subjects, making the dataset applicable for research in CVD risk prediction [18, 17]. Each subject undergoes one to three CT examinations, each producing multiple CT scans employing different CT reconstruction kernels. It has 33,413 CT volumes from 10,395 subjects, and each subject is labeled as either CVD-Positive or CVD-Negative. A subject is considered CVD positive if any cardiovascular abnormalities are reported in their CT screening examinations, or if the subject dies of CVD. CVD-negative subjects have no history of CVD during the clinical trial, no cardiovascular abnormalities reported in any CT scans, and do not die from circulatory system diseases.
@@
We developed the continuous and discrete features joint representation CVD risk prediction system (DeepCVD) in the LDCT-NLST cohort (Fig. 1a). We adopt the three random subsets generated by the prior study [18] entirely, with the training set accounting for 70% (7,268 subjects), the validation set 10% (1,042 subjects), and the independent test set 20% (2,085 subjects). We found that the LDCT-NLST cohort contained some data that did not meet the requirements of our model, such as incomplete field of view (FOV), missing slices, and severe cardiac artifacts. Two radiologists cleansed the dataset and removed the problematic subjects, resulting in 10,325 subjects. The training set comprises 7,227 subjects, the validation set comprises 1,033 subjects, and the test set includes 2,065 subjects. Our goal was to achieve accurate CVD risk prediction using chest CT scans while providing physicians with reliable bases for model decisions.
@@
The proposed DeepCVD was used to identify subjects with CVDs in the LDCT-NLST testing cohort, which consisted of 2,065 patients (583 patients CVD-Positive and 1,482 patients CVD-Negative). The quantitative performance was summarized in Table 1, and the receiver operating characteristic curves (ROCs) of multiple methods were shown in Fig. 1c. DeepCVD achieved an area under the curve (AUC) of 0.875 (95% Confidence Interval (CI), 0.857-0.891), a sensitivity of 0.616 (95% CI, 0.576-0.655), and a specificity of 0.918 (95% CI, 0.903-0.931). From Table 1, it was evident that models using our designed quantitative biomarkers or directly using prior knowledge, such as pericoronary calcification and epicardial fat, whether Xgboost [21], Tri2D-Net [18], or our proposed DeepCVD, demonstrated significantly better specificity compared to models like ResNet34 [22] that did not use prior knowledge. However, the complex encoding of prior knowledge can impair the ability to suppress false positives (non-encoded Xgboost specificity is 0.971, while encoded Tri2D-Net and DeepCVD had specificities of 0.952 and 0.918, respectively). By combining quantitative biomarkers and continuous deep features, DeepCVD increased the sensitivity by 27.0% and the F1 score by 11.9% compared to Tri2D-Net, at the expense of slightly reduced specificity.
To further present the quality of the embeddings of different methods in the LDCT-NLST testing cohort, t-SNE [23] was used to visualize the embeddings of different methods. In Fig. 2, the color of the nodes corresponded to CVD-Positive and CVD-Negative, verifying the discriminative power of the method. As indicated in Fig. 2, DeepCVD and Tri2D-Net can achieve more compact and separated clusters compared with other methods that do not incorporate prior insights.
@@
To assess the generalizability of our proposed DeepCVD, we directly applied the model trained on the LDCT-NLST dataset to the NERC-MBD standard-dose chest CT dataset. DeepCVD achieved an AUC of 0.843 (95% CI, 0.837-0.849), accuracy of 0.819 (95% CI, 0.813-0.824), sensitivity of 0.756 (95% CI, 0.746-0.765), and specificity of 0.870 (95% CI, 0.863-0.876). These results showed significant performance improvement (p ) compared to the previous state-of-the-art approach Tri2D-Net [18] (accuracy increased by 9.8%, sensitivity by 28.4%, F1 score by 17.1%, and AUC by 5.9%, while the specificity remained almost the same). In comparison, other methods had inferior performance that was statistically significant (all p ). Detailed evaluation performance was summarized in Table 2.
@@ Discussion
This study thoroughly validates the effectiveness of DeepCVD. It successfully combines the advantages of discrete and continuous features, improves CVD risk prediction performance, and demonstrates strong generalizability. Multisource feature fusion techniques are also widely used in the medical domain, as feature concatenation is a commonly adopted and effective method for joint feature representation [24]. In the first stage, we concatenate continuous and discrete features in this manner and then feed them into the second stage features joint representation module. We observe an increase in AUC from 0.844 to 0.862 compared to directly using a CVD risk classifier with ResNet34. However, this mode of feature interaction does not fully leverage the respective advantages of quantitative discrete biomarkers and deep continuous features, with specificity increasing from 0.896 to 0.956, while sensitivity decreased from 0.543 to 0.440 (as indicated in Table 1 and Fig. 6c). Furthermore, although our model is trained on LDCT-NLST, it demonstrates remarkable generalization to standard-dose chest CT scans as shown in Table 2. Moreover, our joint representation module outperformed models like Xgboost, even when only discrete features were used, demonstrating the benefits of effective feature interaction for enhancing model performance (shown in Fig. 6d).
@@
We have also discovered that incorporating prior knowledge, either through direct predictions based on CT quantitative biomarkers using Xgboost or by integrating them into deep models such as Tri2D-Net, effectively improves the model’s specificity. Tri2D-Net incorporated pericardial fat and calcification as strong constraints in the prediction model. In the real world, not all CVDs are strongly associated with just two biomarkers, which reduces the model’s sensitivity and its ability to generalize to real-world scenarios. As shown in Fig. 2, although Tri2D-Net also demonstrates high-quality embeddings across the LDCT-NLST test set, its performance significantly drops on the external test set (see Table 2). This is mainly because the external test set includes many subjects with cerebral infarction (refer to Extended Data Fig. 3). Consequently, using CAM [25] for model interpretability also becomes less meaningful. In contrast, our model aligns and flexibly processes discrete and continuous features through the instance-wise feature-gated mechanism to generate more robust embeddings. Then, the soft instance-wise feature interaction mechanism achieves thorough interaction and fusion of features while maintaining relative independence. This ensures that our model optimally balances specificity and sensitivity (as seen in Table 1 and Table 2), and automatically learns the relationships between different CVDs and various biomarkers and deep features (as seen in Fig. 3). This informs physicians about the role of each feature in the decision-making process and opens up possibilities to discover new clinical biomarkers in the CVD domain.
@@
A recent study utilized a mature and fully automated abdominal CT-based algorithm with predefined metrics to quantify aortic calcification, muscle density, the ratio of visceral to subcutaneous fat, liver fat, and bone mineral density for assessing CVD risk [12]. The research demonstrated that the multivariate combination of CT biomarkers could effectively enhance CVD risk prediction performance over traditional risk factors. For example, the combination of four CT-based quantitative biomarkers—aortic calcification, muscle density, the ratio of visceral to subcutaneous fat, and liver fat—resulted in a 2-year AUC of 0.817 (95% CI 0.768-0.866). Although this study used a different retrospective cohort and abdominal non-contrast CT, the results of the two studies both demonstrated the potential CVD risk prediction value of harnessing the rich biometric tissue data embedded within all body CT scans that typically go unused in routine practice. However, this approach requires predefined CT biomarkers, and the combination of these biomarkers to enhance CVD risk prediction performance needs to be validated through repeated experiments. Our approach goes a step further by using deep features to represent those CT biomarkers that cannot or have not yet been predefined, thereby enhancing CVD risk prediction performance. Furthermore, through a unique design, our model can output the contribution of each feature in the decision-making process. This not only provides more insights for doctors but also suppresses those CT biomarkers that do not aid in CVD risk prediction.
@@
In clinical practice, the fusion of imaging information with clinical information can result in increased accuracy, mode informative clinical decision-making, and improved patient outcomes [26]. Our DeepCVD can seamlessly integrate into clinical information. In future work, we aim to expand the discrete features from CT quantitative biomarkers to multimodal biomarkers, incorporating laboratory indicators closely linked to CVDs such as blood pressure, cholesterol, and smoking history into the model to enhance predictive performance. Additionally, we plan to establish connections between biomarkers and specific CVDs through contribution scores, providing actionable guidance for physicians in subsequent diagnoses and achieving greater clinical value.
@@ Methods
This study aims to accurately predict CVD risk using chest CT scans and provide physicians with reliable information for model decisions. Our method is based on two findings and assumptions. First, while well-known quantitative biomarkers from CT scans, such as the coronary artery calcium score, are considered risk factors for CVD [12, 27], we believe that many undefined or difficult-to-quantify biomarkers related to CVD have not been fully exploited within chest CT scans. Fully utilizing these biomarkers can effectively improve the performance of CVD risk prediction. Second, the relevance and contributions of biomarkers should vary across different CVDs, requiring these features to interact sufficiently to address the diversity of CVDs while also maintaining relative independence so that the model can adaptively output the contribution of each biomarker in decision-making.